OpenStudy (anonymous):

What is the sum of a 22–term arithmetic sequence where the first term is 54 and the last term is –30?

OpenStudy (anonymous):

i'll give you a medal!

OpenStudy (nenadmatematika):

Sn=n/2*(a1+an) so I think the sum is \[22/2\times(54-30)=264\]